# Stability of the expanding region of Kerr de Sitter spacetimes

### v1

@@ 2 Covariant ADM formulation of the Einstein equations in parabolic gauge
In this section we decompose a $3+1$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$ with respect to a time-function $s$ which satisfies a parabolic gauge, see (1.12). For a given reference metric $\widetilde{{\bm{g}}}$, this gauge equates the deviation of the corresponding lapse functions to the deviation in the mean curvatures of the leaves $\Sigma_{s}$ of the foliation by level sets of $s$. The Einstein equations then become a first order system for the first and second fundamental forms of $\Sigma_{s}$, coupled to a parabolic equation for the lapse function.
@@ 2.1 Preliminaries of the ADM decomposition
Given a time-function $s$, a time-like vectorfield $T$ is defined by $T^{\mu}=-{\bm{g}}^{\mu\nu}\partial_{\nu}s$. The lapse function $\Phi$ is defined $\Phi^{-2}=-{\bm{g}}(T,T)$, and the unit normal to $\Sigma_{s}$ by $N=\Phi T$. For any choice of coordinates $(x^{1},x^{2},x^{3})$ on $\Sigma_{0}$, we can assign to any point $p\in\Sigma_{s}$ the coordinates $(s,x^{1},x^{2},x^{3})$ if $p=\psi_{s}(q)$, where $\psi_{s}$ is the $1$-parameter group of diffeomorphism generated by $T$, and $q\in\Sigma_{0}$ has coordinates $(x^{1},x^{2},x^{3})$. In these coordinates, the metric takes the form: [^15] As usual, we use the summation convention and Latin indices range in $\{1,2,3\}$, while Greek indices range from 00 to $3$.
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+{g}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}.$ (2.1)
@@
Here $N=\Phi T$ is the unit normal, and we have a coordinate frame $E_{i}=\frac{\partial}{\partial x^{i}}$ which is Lie transported by $T=\frac{\partial}{\partial s}$: $[T,E_{i}]=0$. In the frame $(E_{0}=N,E_{1},E_{2},E_{3})$ the metric components are
${\bm{g}}_{00}=-1\qquad{\bm{g}}_{0i}=0\qquad{\bm{g}}_{ij}=g_{ij}\,.$ (2.3)
@@
and we will express similarly the second variation equation for $\partial_{s}k_{ij}$ as well as the Gauss-Codazzi equations of the embedding of $\Sigma_{s}$ in $\mathcal{M}$ in this frame. The derivations are well-known and can be found for instance in [Chr08].
@@
$\operatorname{Ric}_{ij}=g^{mn}R_{minj}\,.$ (2.6)
Similarly for the Levi-Civita connections of $g,{\bm{g}}$: We denote by $\nabla$ the connection induced by ${\bm{\nabla}}$ on $\Sigma_{s}$.
@@
$\operatorname{Ric}_{ij}[g]+\operatorname{tr}k\>k_{ij}-k^{n}{}_{i}\,k_{nj}={\bm{R}}_{0i0j}+{\bm{\mathrm{Ric}}}_{ij}[{\bm{g}}]={\bm{R}}_{0i0j}+\Lambda g_{ij}$ (2.11)
@@
$\operatorname{tr}_{g}\operatorname{Ric}+(\operatorname{tr}k)^{2}-\lvert k\rvert^{2}=2{\bm{\mathrm{Ric}}}_{00}[{\bm{g}}]+{\bm{\mathrm{tr}}}_{{\bm{g}}}{\bm{\mathrm{Ric}}}=2\Lambda$ (2.13)
@@
these are the constraint equations for the first and second fundamental form, complementing the evolution equations (2.4) for $g$ and (2.12) for $k$.
@@
$R^{a}{}_{bmn}=\partial_{m}\Gamma^{a}_{nb}-\partial_{n}\Gamma^{a}_{mb}+\Gamma^{a}_{mc}\Gamma^{c}_{nb}-\Gamma^{a}_{nc}\Gamma^{c}_{mb}$ (2.16)
@@
$R^{a}{}_{bmn}+R^{a}{}_{mnb}+R^{a}{}_{nbm}=0\,,\quad\text{where }R^{a}{}_{bmn}=g^{ac}R_{cnbm}$ (2.18)
@@ 2.2 System of evolution equations in parabolic gauge
In fact, it is convenient to view the second fundamental form as a $(1,1)$ tensor and work with $k_{i}{}^{j}=g^{cj}k_{ic}$ instead of $k_{ij}$. Then the equations (2.21), (2.23) become:
$\partial_{s}{g}_{ij}=$  $\,2\Phi g_{ja}k_{i}{}^{a}\,,$ (2.24)
$\partial_{s}g^{ij}=$ $-2\Phi g^{ia}k_{a}{}^{j}\,,$ (2.25)
$\partial_{s}k_{i}{}^{j}+\Phi k_{l}{}^{l}k_{i}{}^{j}=$ $\,\nabla_{i}\nabla^{j}\Phi-\Phi\operatorname{Ric}_{i}{}^{j}+\Phi\Lambda\delta_{i}{}^{j}\,.$ (2.26)
In Section 2.5 below, the Ricci curvature $\operatorname{Ric}_{i}{}^{j}$ is suitably expressed in terms of the Christoffel symbols $\Gamma_{ic}^{a}$. We are led to consider, in addition to the (2.24) and (2.26), the following evolution equation for $\Gamma_{s}$:
$\partial_{s}\Gamma_{ic}^{a}=\nabla_{i}(\Phi k_{c}{}^{a})+\nabla_{c}(\Phi k_{i}{}^{a})-g^{ab}g_{cj}\nabla_{b}(\Phi k_{i}{}^{j})$ (2.27)
This is an immediate consequence of (2.24), (2.25) and the formula (2.17):
$\begin{split}\partial_{s}\Gamma^{a}_{ic}=&-2\Phi k_{d}{}^{a}\Gamma^{d}_{ic}+g^{ab}\bigl{(}\partial_{i}(\Phi k_{cb})+\partial_{c}(\Phi k_{ib})-\partial_{b}(\Phi k_{ic})\bigr{)}\\ =&\,g^{ab}\bigl{(}\nabla_{i}(\Phi k_{cb})+\nabla_{c}(\Phi k_{ib})-\nabla_{b}(\Phi k_{ic})\bigr{)}\,.\end{split}$ (2.28)
@@ 2.2.1 Reference metric and gauge
To set up the stability problem, we consider a reference metric:
@@
defined on the same differentiable manifold $\mathcal{M}$, and denote by $\widetilde{{\bm{\mathrm{Ric}}}}_{\mu\nu}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{\mu\nu}$ the components of the Ricci curvature of $\widetilde{{\bm{g}}}$. Also, we denote by ~$\widetilde{\nabla},\widetilde{\Gamma}_{ij}^{a},\widetilde{\operatorname{Ric}}_{ij}=\operatorname{Ric}[\widetilde{g}]_{ij}$ the Levi-Civita connection, Christoffel symbols, and Ricci curvature associated to $\widetilde{g}$.
@@
$\begin{split}\widehat{\Phi}=\Phi-\widetilde{\Phi},\qquad\widehat{g}_{ij}=g_{ij}-\widetilde{g}_{ij},\qquad\widehat{g}^{ij}=g^{ij}-\widetilde{g}^{ij},\\ \widehat{\nabla}=\nabla-\widetilde{\nabla},\qquad\widehat{\Gamma}_{ij}^{a}=\Gamma_{ij}^{a}-\widetilde{\Gamma}_{ij}^{a},\qquad\widehat{k}_{i}{}^{j}=k_{i}{}^{j}-\widetilde{k}_{i}{}^{j},\end{split}$ (2.30)
@@
Remark 2.1.
The main remaining gauge freedom is the choice of the time-function $s$. This is the choice of a lapse function, and in this work we set
$\Phi-\widetilde{\Phi}=k_{l}{}^{l}-\widetilde{k}_{l}{}^{l}$ (2.32)
@@
Remark 2.2.
@@
Remark 2.3.
@@ 2.2.2 First and second variation equations for differences
$\partial_{s}\widehat{g}_{ij}-2H\widehat{g}_{ij}=$  $\,2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}$ (2.37)
$\partial_{s}\widehat{g}^{ij}+2H\widehat{g}^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}-2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\,,$ (2.38)
@@
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\widehat{k}_{i}{}^{j}+\mathfrak{G}_{ic}^{a}$ (2.39)
@@
$+\widehat{k}_{c}{}^{a}\widetilde{\nabla}_{i}\widetilde{\Phi}+\widehat{k}_{i}{}^{a}\widetilde{\nabla}_{c}\widetilde{\Phi}-(\widehat{g}^{ab}g_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widehat{g}_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widetilde{g}_{cj}\widehat{k}_{i}{}^{j})\widetilde{\nabla}_{b}\widetilde{\Phi}$
@@
To derive the equations (2.37), (2.38), (2.39), we use the fact that the corresponding variables of the reference metric satisfy the equations (2.24), (2.25), (2.27), and substract them from the equations satisfied by $g_{ij},g^{ij},\Gamma_{ic}^{a}$. The computations are straightforward. ∎
The following derivation gives the second variation equation for $\widehat{k}_{i}{}^{j}$ and shows that (2.33) is a parabolic gauge. The derivation uses a specific expression for the Ricci curvature, which we present first.
@@
$=$ $\,g^{cj}(\nabla_{a}\Gamma_{ci}^{a}-\nabla_{c}\Gamma_{ia}^{a}+\Gamma^{a}_{cb}\Gamma^{b}_{ai}-\Gamma^{a}_{ab}\Gamma_{ci}^{b}),$
where $\nabla\Gamma$ is interpreted tensorially, e.g.,
Alternatively, we write using the pair symmetry of the curvature tensor
@@
The motivation for these manipulations will be discussed in Remark 2.8. We now return to (2.26).
@@
$\partial_{s}\widehat{\Phi}-\Delta_{g}\widehat{\Phi}+2H\widehat{\Phi}=\mathfrak{F}+\widetilde{\mathfrak{I}}_{\Phi}$ (2.45)
@@
$+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j})+\frac{2}{3}(\widehat{\Phi}g^{ab}+\widetilde{\Phi}\widehat{g}^{ab})(\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j})$
$\mathfrak{F}=$ $-2\widehat{\Phi}\widehat{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Phi\widehat{k}_{i}{}^{j}\widehat{k}_{j}{}^{i}-\widehat{\Phi}(\widetilde{k}_{i}{}^{j}\widetilde{k}_{j}{}^{i}-\Lambda)-\widetilde{\Phi}\widehat{k}_{i}{}^{j}(\widetilde{k}_{j}{}^{i}-H\widetilde{\Phi}^{-1}\delta_{j}{}^{i})$ (2.47)
$+\widehat{g}^{ab}\partial_{a}\partial_{b}\widetilde{\Phi}-g^{ab}\widehat{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}-\widehat{g}^{ab}\widetilde{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}$
The terms (~$(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j},\widetilde{\mathfrak{I}}_{\Phi}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
(~$(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}=-\widetilde{\Phi}\Bigl{(}\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}\Bigr{)},\qquad\widetilde{\mathfrak{J}}_{\Phi}=\widetilde{\Phi}\Bigl{(}\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\Lambda\Bigr{)}\,.$ (2.48)
@@
Subtracting it from (2.26) and using (2.33) then results in

### v2

@@ Acknowledgements.
Recent developments.
Since the first appearance of this preprint, Hintz and Vasy have given an alternative proof of the stability of the cosmological region in generalised harmonic gauge [HV24], allowing in particular a localisation near timelike infinity.
@@ 2 Covariant ADM formulation of the Einstein equations in parabolic gauge
In this section we derive the Einstein equations for a decomposition of a $1+3$-dimensional Lorentzian manifold $(\mathcal{M},{\bm{g}})$ with respect to a time-function $s$ which satisfies a parabolic gauge, see (1.13). For a given reference metric $\widetilde{{\bm{g}}}$, this gauge equates the deviation of the corresponding lapse functions to the deviation in the mean curvatures of the leaves $\Sigma_{s}$ of the foliation. The Einstein equations then become a first order system for the first and second fundamental forms of $\Sigma_{s}$, coupled to a parabolic equation for the lapse function.
@@ 2.1 Preliminaries of the ADM decomposition
We recall various geometric notions in the ADM formalism, all defined locally with respect to a differentiable time-function $s$; for an introduction to this formulation see for instance [Chr08].
Given any time-function $s$ (namely a differentiable function $s$ on $\mathcal{M}$ with the property that $V(s)>0$ for any future-directed timelike vectorfield $V$), we denote by $\Sigma_{s}$ the level sets of $s$ (a family of spacelike hypersurfaces in $\mathcal{M}$), and by $T$ the timelike vectorfield colinear to the normal $N$ defined by $T^{\mu}=-{\bm{g}}^{\mu\nu}\partial_{\nu}s$. The lapse function $\Phi$ is defined $\Phi^{-2}=-{\bm{g}}(T,T)$, and the unit normal to $\Sigma_{s}$ is then given by $N=\Phi T$. For any choice of coordinates $(x^{1},x^{2},x^{3})$ on a given level set $\Sigma_{s_{0}}$, we can assign to any point $p\in\Sigma_{s}$ the coordinates $(s,x^{1},x^{2},x^{3})$ if $p=\psi_{t}(q)$, $t=s-s_{0}$, where $\psi_{t}$ is the $1$-parameter group of diffeomorphism generated by $T$, and $q\in\Sigma_{s_{0}}$ has coordinates $(x^{1},x^{2},x^{3})$. In these coordinates, the metric takes the form: [^16] As usual, we use the summation convention and Latin indices range in $\{1,2,3\}$, while Greek indices range in $\{0,1,2,3\}$.
${\bm{g}}=-\Phi^{2}\mathrm{d}s^{2}+{g}_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (2.1)
@@
Here $N=\Phi T$ is the unit normal, and we have a coordinate frame $E_{i}=\frac{\partial}{\partial x^{i}}$ which is Lie transported by $T=\frac{\partial}{\partial s}$: $T,E_{i}]=0$. In the frame $(E_{0}=N,E_{1},E_{2},E_{3})$ the metric components are
${\bm{g}}_{00}=-1\,,\qquad{\bm{g}}_{0i}=0\,,\qquad{\bm{g}}_{ij}=g_{ij}\,.$ (2.3)
@@
and we will express similarly the second variation equation for $\partial_{s}k_{ij}$ as well as the Gauss-Codazzi equations of the embedding of $\Sigma_{s}$ in $\mathcal{M}$ in this frame.
@@
$\operatorname{Ric}_{ij}=g^{mn}R_{minj}$ (2.6)
[…]
@@
$\operatorname{Ric}_{ij}[g]+\operatorname{tr}k\>k_{ij}-k^{n}{}_{i}\,k_{nj}={\bm{R}}_{0i0j}+{\bm{\mathrm{Ric}}}_{ij}[{\bm{g}}]={\bm{R}}_{0i0j}+\Lambda g_{ij}\,$, (2.11)
@@
$\operatorname{tr}_{g}\operatorname{Ric}+(\operatorname{tr}k)^{2}-\lvert k\rvert^{2}=2{\bm{\mathrm{Ric}}}_{00}[{\bm{g}}]+{\bm{\mathrm{tr}}}_{{\bm{g}}}{\bm{\mathrm{Ric}}}=2\Lambda\,$, (2.13)
@@
these are the constraint equations for the first and second fundamental form, complementing the evolution equations (2.4) for $g$ and (2.12) for $k$. Combining (2.7), (2.13), we also have the following evolution equation for the mean curvature:
$\partial_{s}\mathrm{tr}k=\Delta_{g}\Phi-\Phi|k|^{2}-\Phi{\bf Ric}_{00}[{\bm{g}}].$ (2.16)
@@
$R^{a}{}_{bmn}=\partial_{m}\Gamma^{a}_{nb}-\partial_{n}\Gamma^{a}_{mb}+\Gamma^{a}_{mc}\Gamma^{c}_{nb}-\Gamma^{a}_{nc}\Gamma^{c}_{mb}\,$, (2.17)
@@
$R^{a}{}_{bmn}+R^{a}{}_{mnb}+R^{a}{}_{nbm}=0\,,\quad\text{where }R^{a}{}_{bmn}=g^{ac}R_{cnbm}\,$, (2.19)
@@ 2.2 System of evolution equations in parabolic gauge
In fact, it is convenient to view the second fundamental form as a $(1,1)$ tensor and work with $k_{i}{}^{j}=g^{cj}k_{ic}$ instead of $k_{ij}$. Then the equations (2.22a), (2.23) become:
$\partial_{s}{g}_{ij}=$  $\,2\Phi g_{ja}k_{i}{}^{a}$ (2.24a)
$\partial_{s}g^{ij}=$ $-2\Phi g^{ia}k_{a}{}^{j}$ (2.24b)
$\partial_{s}k_{i}{}^{j}+\Phi k_{l}{}^{l}k_{i}{}^{j}$= $\,\nabla_{i}\nabla^{j}\Phi-\Phi\operatorname{Ric}_{i}{}^{j}+\Phi\Lambda\delta_{i}{}^{j}$ (2.24c)
In Section 2.6 below, the Ricci curvature $\operatorname{Ric}_{i}{}^{j}$ is suitably expressed in terms of the Christoffel symbols $\Gamma_{ic}^{a}$. We are led to consider, in addition to the (2.24a) and (2.24c), the following evolution equation for $\Gamma_{s}$:
[…]
This is an immediate consequence of (2.24a), (2.24b) and the formula (2.18):
$\begin{split}\partial_{s}\Gamma^{a}_{ic}=&-2\Phi k_{d}{}^{a}\Gamma^{d}_{ic}+g^{ab}\bigl(\partial_{i}(\Phi k_{cb})+\partial_{c}(\Phi k_{ib})-\partial_{b}(\Phi k_{ic})\bigr)\\ =&\,g^{ab}\bigl(\nabla_{i}(\Phi k_{cb})+\nabla_{c}(\Phi k_{ib})-\nabla_{b}(\Phi k_{ic})\bigr)\end{split}$ (2.26)
@@ 2.2.1 Reference metric and gauge
To set up the stability problem, we will make the choice of a reference metric in Section 3:
@@
defined on the same differentiable manifold $\mathcal{M}$, and denote by $\widetilde{{\bm{\mathrm{Ric}}}}_{\mu\nu}={\bm{\mathrm{Ric}}}[\widetilde{{\bm{g}}}]_{\mu\nu}$ the components of the Ricci curvature of $\widetilde{{\bm{g}}}$. Also, we denote by $\widetilde{\nabla},\widetilde{\Gamma}_{ij}^{a},\widetilde{\operatorname{Ric}}_{ij}=\operatorname{Ric}[\widetilde{g}]_{ij}$ the Levi-Civita connection, Christoffel symbols, and Ricci curvature associated to $\widetilde{g}$.
@@
[…]
@@
The remaining gauge freedom is the choice of the time-function $s$. This is the choice of a lapse function, and in this work we set
$\Phi-\widetilde{\Phi}=k_{l}{}^{l}-\widetilde{k}_{l}{}^{l}\,$, (2.30)
@@
Remark 2.1 (Geometric interpretation of differences).
In the definition of the hatted quantities (2.28), a common identification of the coordinate vectorfields is made. Strictly speaking, $\bm{g}$} and $\widetilde{{\bm{g}}}$ are metrics on different manifolds. However, they are expressed in the same chart. If we view $(s,x^{1}=t,x^{2},x^{3})$ as standard coordinates for $\widetilde{{\bm{g}}}$, then $\widetilde{\Phi}^{2}$, and $\widetilde{g}_{ij}$ refer to the components of $\widetilde{{\bm{g}}}$ with respect to the coordinate basis $(\partial_{s},\partial_{x^{i}})$ in the chart. However, $\Phi^{2}$, and $g_{ij}$ are the components of the metric $\bm{g}$} with respect to the basis $(\mathrm{d}\psi\cdot\partial_{s},\mathrm{d}\psi\cdot\partial_{x^{i}})$, where $\psi:(s,x)\mapsto\psi(s,x)\in\mathcal{R}$ is a diffeomorphism, a chart for $(\mathcal{R},{\bm{g}})$. These are the coordinate vectorfields in the given chart and typically also denoted by $(\partial_{s},\partial_{x^{i}})$. This allows us to view $g_{ij}$ and $\widetilde{g}_{ij}$ as functions of $(s,x^{1},x^{2},x^{3})$ in the same chart. In other words, the hatted quantities can be viewed as the components of $\psi^{\ast}{\bm{g}}-\widetilde{{\bm{g}}}$ is the standard coordinate chart of $\widetilde{{\bm{g}}}$. In particular, with this identification both $g_{ij}$ and $\widetilde{g}_{ij}$ are $\Sigma_{s}$-tensors.
Remark 2.2 (Contractions with the metric).
Remark 2.3 (Elliptic gauge choices).
@@
Remark 2.4 (Identity for differences of Christoffel symbols).
@@ 2.2.2 First and second variation equations for differences
$\partial_{s}\widehat{g}_{ij}-2H\widehat{g}_{ij}=$  $\,2\Phi g_{ja}\widehat{k}_{i}{}^{a}+2\widehat{\Phi}g_{ja}\widetilde{k}_{i}{}^{a}+2\widetilde{\Phi}(\widetilde{k}_{i}{}^{a}-H\widetilde{\Phi}^{-1}\delta_{i}{}^{a})\widehat{g}_{ja}\,$, (2.35a)
$\partial_{s}\widehat{g}^{ij}+2H\widehat{g}^{ij}=$ $-2\Phi g^{ia}\widehat{k}_{a}{}^{j}-2\widehat{\Phi}g^{ia}\widetilde{k}_{a}{}^{j}-2\widetilde{\Phi}(\widetilde{k}_{a}{}^{j}-H\widetilde{\Phi}^{-1}\delta_{a}{}^{j})\widehat{g}^{ia}\,$, (2.35b)
@@
$\partial_{s}\widehat{\Gamma}_{ic}^{a}=$ $\,\Phi\nabla_{i}\widehat{k}_{c}{}^{a}+\Phi\nabla_{c}\widehat{k}_{i}{}^{a}-g^{ab}g_{cj}\Phi\nabla_{b}\widehat{k}_{i}{}^{j}+\mathfrak{G}_{ic}^{a}\,$, (2.36)
@@
$+\widehat{k}_{c}{}^{a}\widetilde{\nabla}_{i}\widetilde{\Phi}+\widehat{k}_{i}{}^{a}\widetilde{\nabla}_{c}\widetilde{\Phi}-(\widehat{g}^{ab}g_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widehat{g}_{cj}k_{i}{}^{j}+\widetilde{g}^{ab}\widetilde{g}_{cj}\widehat{k}_{i}{}^{j})\widetilde{\nabla}_{b}\widetilde{\Phi}\,$.
@@
To derive the equations (2.35a), (2.35b), (2.36), we use the fact that the corresponding variables of the reference metric satisfy the equations (2.24a), (2.24b), (2.25), and subtract them from the equations satisfied by $g_{ij},g^{ij},\Gamma_{ic}^{a}$. The computations are straightforward. ∎
While the first variation formulas do not rely on the gauge condition, it is used in the following derivation of the second variation equation for $\widehat{k}_{i}{}^{j}$ which shows that (2.31) is a parabolic gauge. The derivation uses a specific expression for the Ricci curvature, which we present first.
@@
$=$ $\,g^{cj}(\nabla_{a}\Gamma_{ci}^{a}-\nabla_{c}\Gamma_{ia}^{a}+\Gamma^{a}_{cb}\Gamma^{b}_{ai}-\Gamma^{a}_{ab}\Gamma_{ci}^{b})$
Alternatively, we write using the pair symmetry of the curvature tensor:
@@
The motivation for these manipulations will be discussed in Remark 2.9. In fact, we have already seen in (2.36) that the evolution equation for $\widehat{\Gamma}$ contains derivatives of $\widehat{k}$. Now returning to (2.24c), we will see that the evolution equation for $\widehat{k}$ can be expressed in terms of derivatives of $\widehat{\Gamma}$, in such a way that in the derivation of the energy estimate, which couples these evolution equations, several terms cancel. This uses the formulas for the Ricci curvature in Lemma 2.6.
@@
$+\frac{1}{3}\Phi g^{cj}(\nabla_{c}\widehat{\Gamma}_{ia}^{a}-\nabla_{a}\widehat{\Gamma}_{ci}^{a})+\frac{2}{3}\Phi g^{ab}(\nabla_{a}\widehat{\Gamma}^{j}_{bi}-\nabla_{i}\widehat{\Gamma}^{j}_{ab})\,$,
@@
$\partial_{s}\widehat{\Phi}-\Delta_{g}\widehat{\Phi}+2H\widehat{\Phi}=\mathfrak{F}+\widetilde{\mathfrak{I}}_{\Phi}\,$, (2.42)
@@
$+\frac{2}{3}\Phi g^{ab}(\widetilde{\Gamma}^{c}_{ab}\widehat{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widehat{\Gamma}_{ac}^{j})+\frac{2}{3}(\widehat{\Phi}g^{ab}+\widetilde{\Phi}\widehat{g}^{ab})(\widetilde{\Gamma}^{c}_{ab}\widetilde{\Gamma}^{j}_{ci}-\widetilde{\Gamma}^{c}_{ib}\widetilde{\Gamma}_{ac}^{j})\,$,
[…]
$+\widehat{g}^{ab}\partial_{a}\partial_{b}\widetilde{\Phi}-g^{ab}\widehat{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}-\widehat{g}^{ab}\widetilde{\Gamma}_{ab}^{c}\partial_{c}\widetilde{\Phi}\,$.
The terms $(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j},\widetilde{\mathfrak{I}}_{\Phi}$ only contain variables of the reference metric $\widetilde{{\bm{g}}}$ and are equal to:
$(\widetilde{\mathfrak{I}}_{k})_{i}{}^{j}=-\widetilde{\Phi}\Bigl(\widetilde{{\bm{\mathrm{Ric}}}}_{i}{}^{j}-\Lambda\delta_{i}{}^{j}\Bigr),\qquad\widetilde{\mathfrak{I}}_{\Phi}=\widetilde{\Phi}\Bigl(\widetilde{{\bm{\mathrm{Ric}}}}_{00}+\Lambda\Bigr)\,.$ (2.45)
@@
Subtracting it from (2.24c) and using (2.31) then results in